# Exploring the Potential of Large Language Models for Improving Digital Forensic Investigation Efficiency

### v2

@@ 3.6 Retrieval-augmented Generation
A major advantage of RAG is its ability to reduce the hallucination problem in LLMs, enabling them to generate more accurate and current information [131, 132]. The architecture of RAG consists of a knowledge base and a retriever model. The retriever model converts input prompts and content from the knowledge base into vectors. The user prompts are then appended with the most relevant content from the knowledge base, and this augmented prompt is sent to the base model to generate a more accurate response [132].
@@ 3.7 Limitations and Risks
As explored in the preceding sections, LLMs appear to possess a vast range of capabilities. However, it is crucial to acknowledge that they are not without limitations and risks. In multimodal LLMs, it is a common issue that they are over-reliant [133]. Hadi et al. [134] highlighted significant drawbacks associated with LLMs, including issues such as bias, explainability challenges, reasoning errors, logical errors, hallucinations, vulnerability to prompt injections, and spelling and grammar errors. These limitations underscore the importance of a cautious and critical approach when utilising LLMs in various applications. Additionally, the literature shows limitations in LLMs, including statistical inconsistency, the absence of emotional attributes in linguistic responses, and challenges related to fact verification [135, 136]. These factors contribute to a comprehensive understanding of the constraints and potential shortcomings when working with LLMs. Thapa et al. [137] contend that while LLMs can indeed diminish the time and costs associated with annotation tasks, they fall short of completely supplanting human annotation. This is because they struggle with intricate linguistic constructions, such as idioms, irony, sarcasm, and metaphor, which can potentially impact the precision of annotations.
@@
Similar limitations are associated with MLLMs, such as over-reliance on training data, sensitivity to word order in the input prompts, and vulnerability to prompts containing extra knowledge [138]. There are several more concerns associated with LLMs including restricted input and output text lengths, limited comprehension of syntax, ethical considerations with the generated information, constraints with multilingual capabilities, elevated costs associated with training and maintenance, inadequate understanding of human behaviours, and limited ability to learn incrementally [134, 139, 140].
@@
Despite their considerable capabilities, LLMs are not without risks. Bommasani et al. [141], Lund and Wang [142], Rahman and Santacana [143] provide comprehensive overviews of risks linked to LLMs. These include the homogenisation of outcomes, wherein defects or biases from the foundation model are inherited by all downstream models. There is also the risk of a monopoly for the owners of the foundation models, leading to increased centralisation of power in a single entity. Ethical and legal concerns are intertwined with privacy and intellectual property issues. Additionally, there are economic and environmental impacts, raising concerns about the potential displacement of human workers. Furthermore, inequity and misuse of LLMs, such as the creation of deepfakes and their application in criminal and unethical activities, pose additional challenges. Given that LLMs do not inherently prioritise the accuracy of information, Bender et al. [144] have underscored the risk of generating social turbulence, especially, when employed on social media platforms. Additionally, the use of LLMs is associated with significant costs, leading to a direct environmental impact due to their substantial energy consumption [145].
@@
In response to the increasing prominence of LLMs and Artificial General Intelligence (AGI), the EU Artificial Intelligence Act (AIA) presently classifies AI systems into four risk categories based on broad fields of application. However, this categorisation approach may result in inaccurate risk estimation and ineffective enforcement. To rectify this issue, the authors advocate for a risk assessment model that aligns the risk categories with specific AI scenarios. This model integrates the AIA with the risk assessment approach employed by the Intergovernmental Panel on Climate Change (IPCC) and relevant literature, facilitating a more precise estimation of the magnitude of AI-related risks [147].
The Beyond the Imitation Game benchmark (BIG-bench), serves as an evaluation framework for LLMs. It encompasses 204 distinct language-related tasks. These span contextual and context-free question-answering, reading comprehension, logical reasoning, etc. [148]. It is acknowledged that the challenge of social biases and English language dependency persists across nearly all LLMs.
When employing LLM-based agents, it is imperative to address challenges associated with LLM-based multi-agent frameworks as well. Handling many defined agents may necessitate substantial computational resources and memory, thereby mandating high-end computing infrastructure for seamless operations. Furthermore, the absence of a standardised comprehensive benchmarking system to evaluate the behaviour of such agents underscores the limitations inherent in the development of LLM-based multi-agent systems. These challenges underscore the need for further research and refinement in this domain to enhance the efficiency and effectiveness of LLM-based multi-agent frameworks [149].
Similar to these risks, the Open Web Application Security Project (OWASP) has identified ten major risk factors related to LLMs. These risks include prompt injection, insecure output handling, training data poisoning, model denial of service, supply chain vulnerabilities, sensitive information disclosure, insecure plugins, excessive agency, overreliance, and model theft. Despite these threat factors, OWASP also emphasised the need for regulatory bodies to oversee LLMs across various domains and recommended the implementation of risk management programs that incorporate the checklist provided by OWASP [^2] https://owasp.org/www-project-top-10-for-large-language-model-applications.
@@ 4 Large Language Models For Digital Forensics
Section 4 summarises existing work with LLMs in DF, the feasibility of employing them, and potential future directions. As discussed in Sections 2 and 3, despite the widespread use of LLMs in various fields to enhance the efficiency and accuracy of tasks within specific domains, their application in the field of DF is still relatively new.
Conducting a thorough analysis of the utilisation of LLMs in conjunction with the stages of the DF process model, as highlighted in Section 1.1.1, proves to be a valuable undertaking.
@@ 4.1 Incident Recognition Phase
In the initial phase of Casey’s DF process model, which delineates the recognition of an incident, LLMs can serve as a valuable detection mechanism [150]. In cybercrime cases, the primary artefacts often involve data logs, data dumps and network dumps. Fine-tuning an LLM to monitor text-based logs and related files enables it to discern and identify potential or ongoing incidents within the environment. In network-related activities, anomaly detection plays a pivotal role in initiating an incident response. Various existing anomaly detection techniques are employed in systems for this purpose. Leveraging its capability to identify patterns in a series of text data sets, LLMs exhibit potential as an Intrusion Detection System (IDS) within such systems [151].
@@ 4.2 Collection Phase
While evidence collection or seizure traditionally involves physical tasks requiring human interaction, LLMs can play a role in identifying and listing potential pieces of evidence at a crime scene. For instance, in the examination of photographs or video records from a crime scene, an investigator can enlist the assistance of a MLLM like LLaVa, GPT-4, or VisionLLM. These models are capable of processing information within the images and generating a text-based output, facilitating the interpretation and categorisation of visual data. While this task may seem simple and within the capabilities of a human agent, the efficiency becomes particularly evident, when dealing with a massive-scale investigation involving thousands of collected artefacts and photographs. Utilising an MLLM for initial processing can significantly save time, with human agents then focusing on the crucial task of verification and validation.
@@ 4.3 Preservation/Acquisition Phase
Preserving evidence is centred on maintaining integrity. To achieve this, various tools such as EnCase and FTK Imager have been employed, aiding investigators in streamlining their work processes [152]. In the context of preserving disk evidence, it becomes feasible for an investigator to articulate their requirements in natural language, LLM to generate source code tailored to the specific needs. Indeed, LLMs specialised in code generation, such as StarCoder, Code LLaMA, and others, can be fine-tuned and retrained for specific tasks, including those related to preserving disk evidence through customised code generation. In certain instances, gathering live data for forensic investigations becomes crucial, particularly data collected at the crime scene. For this purpose, investigators may leverage DFaaS platforms like Hansken. Hansken possesses the capability to amalgamate custom extraction APIs for data extractions, and these APIs can be developed using code-generative LLMs [29]. This approach enhances the adaptability and efficiency of the investigative process.
As stated in Section 3.5, the automation of code generation and unit testing can be facilitated by autonomous agents that utilise LLMs as their core. AutoGen, being an open-source framework, provides the means to develop AI agents tailored for specific tasks. These AutoGen agents are not only customizable and conversational but can also operate in diverse modes, employing combinations of LLMs, human inputs, and various tools [127]. Positively, automated agents, particularly those developed within frameworks like AutoGen, can be effectively employed in the preservation phase. Their capabilities will enable the automated preservation within this phase, streamlining and enhancing efficiency in the preservation of digital evidence [127].
@@ 4.4 Examination Phase
This phase constitutes a pivotal component of the investigation, playing a crucial role in elucidating the case through activities such as data recovery, collection, reduction, and classification. For each of these components, LLMs fine-tuned for scripting can significantly assist, especially at a larger scale. Within these components, tasks such as keyword search, file recovery, pattern matching, and fragment reassembly can be achieved with minimal technical knowledge using LLMs. LLMs can provide valuable assistance in these tasks by generating new codes, crafting regular expressions, generating passwords and/or password hash lists for decryption, and creating sample logs or files. LLMs can generate a set of instructions, queries, and Application Programming Interface (API) validations from natural language provided by a human. This opens up the possibility of integrating third-party tools like Scapy, tshark, John the Ripper, and others seamlessly into the investigative process, enhancing the toolkit available for DF investigations and the ability to automate these processes enhances efficiency and effectiveness in the examination phase of the investigation.
The use of LAMs and VoT techniques in the examination phase can significantly enhance the efficiency of an investigation. Since LAMs and VoT specifically focus on task manipulation, investigators can offload some examination work to an LAM, which will then generate the final results from a series of subtasks. This approach can allow investigators to focus on higher-level analysis and decision-making, thereby streamlining the overall investigative process.
@@ 4.5 Analysis Phase
The analysis phase involves comprehending the incident and arriving at a conclusive understanding based on the information gathered during the examination phase. As also highlighted in Section 1.1.3, it has been demonstrated that LLMs are effective in case analysis [30]. The use of MLLMs, which possess the capability to interpret images, broadens the scope for analysing a crime case more comprehensively. LLMs can be specifically fine-tuned for the analysis of various data types, including log files, email contents, chat transcripts, call records, file metadata, hex dumps, memory dumps, and registry hives. Incorporating contents such as event logs, timestamps, and network traffic captures further enables the effective recreation of incidents by correlating each data set with the assistance of LLMs. Also, audio and video-specific MLLMs can assist in analysing the contents within these formats. This specialised capability can significantly reduce the time investigators spend analysing audio and video data during investigations.
The use of automated agents can distribute the analysis workload effectively. Simultaneously, leveraging Augmented Large Language Models (ALLMs) and RAG techniques can enhance real-time knowledge retrieval, thereby improving the accuracy of analysis and decision-making processes [153]. For example, integrating a source of intelligence with an RAG system can assist investigators in connecting the dots during a DF investigation.
@@ 4.6 Reporting Phase
The quality and validity of evidence, along with the thoroughness of the analysis, are encapsulated in the final report. The reporting phase holds significant weight, as the entire judgement may hinge on this crucial stage. Notably, DF is experiencing heightened scrutiny about the quality of the reports, emphasising the importance of precision and clarity in this phase [154]. As pointed out by Champod et al. [155], there is no standard framework for evaluating and reporting scientific findings to authorities and stakeholders. To provide assistance and alleviate scrutiny, incorporating LLMs for report creation is a viable solution. While by definition, LLMs do not produce deterministic outputs, they could be retrained using standards such as ISO/IEC 27043:2015, the international standard for the DF investigation process [156]. A first feasibility study has already been released by Michelet and Breitinger [157]. Naturally, LLMs could also help to automate forensic reports and alternative formats (i.e., not MS Word) such as HTML or Latex.
4.7 Other Possibilities
In general, LLMs can be effectively used in each phase of the DF process model. By integrating the life cycle of the DF process with automated agents, it becomes feasible to establish a fully automated DF process throughout an investigation. For instance, if automated agents are assigned specific roles and maintain synchronisation within each agent, employing frameworks such as AutoGen, there is a substantial potential to expedite the completion of one or multiple DF cases compared to relying solely on human agents. This approach improves efficiency, speed, and consistency in the DF investigation process.
Scanlon et al. [13] highlights that LLMs can play an important role in DF teaching scenarios. This involvement extends to activities such as storyboarding, synthetic content creation, and synthetic character profiling. Additionally, fine-tuned LLMs prove to be valuable tools for establishing evaluation criteria for DF agents, facilitating the creation of exams, measuring the accuracy and efficiency of investigations through an automated weighing system, and effectively presenting a case to non-technical or non-domain individuals, such as judges.
Additionally, to provide a comprehensive understanding of the potential utilisation of LLMs in DF, Table LABEL:DF_example_prompts elucidates sample DF functionalities within the framework of the National Institute of Standards and Technology (NIST) Computer Forensics Tool Testing Program (CFTT), highlighting the usability of LLMs and example prompts. The CFTT project establishes overarching specifications to assess the capabilities of DF tools, a framework adopted by numerous prominent free and commercial DF tools  [^3] https://www.nist.gov/itl/ssd/software-quality-group/computer-forensics-tool-testing-program-cftt.
@@ 4.7 Other Possibilities
Also, to explore the applicability of LLMs to DF, a simple matrix was defined in Table LABEL:potential_LLMsin_DFphases. This matrix showcases the high, medium, and low potential DF phases where LLMs can be utilised. The usability rankings are based on factors such as the extent of human intervention and effort required, the prospect of automation, and the applicability of human expert knowledge.
{longtblr}
[ caption = Potential of LLMs in Digital Forensic Phases, label = potential_LLMsin_DFphases, ] width = colspec = Q[130]Q[490]Q[170], cells = c, hlines, vlines, Potential & Description DF Phases
High Demands minimal human expert knowledge. Exclusively deals with digital evidence. Fully amenable to automation. Examination, Analysis, Reporting
Medium Necessitates human expert knowledge, though it is not a critical component. Focuses on handling digital evidence, with minimal involvement of physical evidence. Automation is feasible and can be implemented. Preservation
Low Requires a high level of human expert knowledge. Often requires identifying/handling physical evidence. Minimal or no scope for automation. Collection

### v3

@@ 3.6 Retrieval-augmented Generation
A major advantage of RAG is its ability to reduce the hallucination problem in LLMs, allowing them to generate more accurate and current information [134, 135]. The architecture of RAG consists of a knowledge base and a retriever model. The retriever model converts input prompts and content from the knowledge base to vectors. The user prompts are then appended with the most relevant content from the knowledge base, and this augmented prompt is sent to the base model to generate a more accurate response [135].
@@ 3.7 Limitations and Risks
As explored in the preceding sections, LLMs appear to possess a vast range of capabilities. However, it is crucial to acknowledge that they are not without limitations and risks. In multimodal LLMs, it is a common problem that they are over-reliant [136]. There are also potentially significant drawbacks associated with LLMs, including issues such as bias, explainability challenges, reasoning errors, logical errors, hallucinations, vulnerability to prompt injections, and spelling and grammar errors. These limitations underscore the importance of a cautious and critical approach when using LLMs in various applications. Furthermore, the literature shows limitations in LLMs, including statistical inconsistency, the absence of emotional attributes in linguistic responses, and challenges related to fact verification [137, 138]. These factors contribute to a comprehensive understanding of the constraints and potential shortcomings when working with LLMs. Thapa et al. [139] contend that while LLMs can indeed reduce the time and costs associated with annotation tasks, they are not completely supplanting human annotation. This is because they struggle with intricate linguistic constructions, such as idioms, irony, sarcasm, and metaphor, which can potentially impact the precision of annotations.
@@
Similar limitations are associated with MLLMs, such as over-reliance on training data, sensitivity to word order in input prompts, and vulnerability to prompts containing additional knowledge [140]. There are several more concerns associated with LLMs, including restricted text input and output lengths, limited comprehension of syntax, ethical considerations with the generated information, constraints with multilingual capabilities, elevated costs associated with training and maintenance, inadequate understanding of human behaviours and limited ability to learn incrementally [141, 142].
@@
Despite their considerable capabilities, LLMs are not without risks. Wiggins and Tejani [143], Lund and Wang [144], Rahman and Santacana [145] provide comprehensive overviews of risks linked to LLMs. These include the homogenisation of results, whereby defects or biases from the foundation model are inherited by all downstream models. There is also the risk of monopolistic control by foundation model owners, potentially concentrating decision-making power, resource access, and influence over model usage in a single entity. Ethical and legal concerns are intertwined with concerns about privacy and intellectual property. In addition, there are economic and environmental impacts that raise concerns about the potential displacement of human workers. Furthermore, inequity and misuse of LLMs, such as the creation of deepfakes and their application in criminal and unethical activities, pose additional challenges. Given that LLMs do not inherently prioritise the precision of information, Bender et al. [146] have highlighted the risk of generating social turbulence, especially when used on social media platforms. Furthermore, the use of LLMs is associated with significant costs, leading to a direct environmental impact due to their substantial energy consumption [147].
@@
The Beyond the Imitation Game benchmark (BIG-bench), serves as an evaluation framework for LLMs. It encompasses 204 distinct language-related tasks. These span contextual and context-free question-answering, reading comprehension, logical reasoning, etc. [149]. It is acknowledged that the challenge of social biases and dependency on the English language persists in almost all LLMs.
When employing LLM-based agents, it is imperative to address challenges associated with LLM-based multi-agent frameworks as well. The handling of many defined agents may necessitate substantial computational resources and memory, thus mandating high-end computing infrastructure for seamless operations. Furthermore, the absence of a standardised comprehensive benchmarking system to evaluate the behaviour of such agents underscores the limitations inherent in the development of LLM-based multi-agent systems. These challenges underscore the need for further research and refinement in this domain to enhance the efficiency and effectiveness of LLM-based multi-agent frameworks [150].
Similarly to these risks, the Open Web Application Security Project (OWASP) has identified ten major risk factors related to LLMs. These risks include training data poisoning, prompt injection, denial of service, insecure output handling, supply chain vulnerabilities, sensitive information leakage, excessive agency, insecure plugins, overreliance, and model theft of data. Despite these threat factors, OWASP also stressed the need for regulatory bodies to supervise LLMs in various domains and recommended the implementation of risk management programmes that incorporate the checklist provided by OWASP [^4] https://owasp.org/www-project-top-10-for-large-language-model-applications.
The EU’s Artificial Intelligence Act (AIA) proposes a framework for categorising AI applications based on their associated risk levels, with the primary aim of safeguarding human rights and maintaining ethical standards in AI deployment [151]. Within the AIA, AI applications are divided into categories such as “unacceptable risk”, which includes practices such as exploiting vulnerabilities and social-ranking techniques due to their potential for individual manipulation and impact on fundamental rights. These categories have relevance to DF, where the ethical application of LLMs must balance the advantages of automation with the imperative to address privacy concerns. Since DF involves sensitive data and influences legal outcomes, the use of LLMs must align with AIA’s risk-based principles, ensuring transparency, accountability, and fair application. In future AI applications within DF, it is essential to implement appropriate measures to prevent biases and hallucinations to mitigate the risk of misuse of AI.
@@ 4 Large Language Models For Digital Forensics
Section 4 summarises existing work with LLMs in DF, the feasibility of employing them, and potential future directions. As discussed in Sections 2 and 3, despite the widespread use of LLMs in various fields to improve the efficiency and accuracy of tasks within specific domains, their application in the field of DF is still relatively new.
Conducting a thorough analysis of LLMs use in conjunction with the stages of the DF process model, as highlighted in Section 1.1.1, proves to be a valuable undertaking.
@@ 4.1 Incident Recognition Phase
In the initial phase of Casey’s DF process model, which delineates the recognition of an incident, LLMs can serve as a valuable detection mechanism [152]. In cybercrime cases, the primary artefacts often involve data logs, data dumps and network dumps. Fine-tuning an LLM to monitor text-based logs and related files enables it to discern and identify potential or ongoing incidents within the environment. In network-related activities, anomaly detection plays a pivotal role in initiating an incident response. Various existing anomaly detection techniques are employed in systems for this purpose. Using their ability to identify patterns in a series of text data sets, LLMs exhibit potential as an Intrusion Detection System (IDS) within such systems [33, 153].
For instance, Kan et al. [154] introduces Mobile-LLaMA, an open source mobile network-specialised LLM, fine-tuned through instructional data to enhance their capabilities for network analysis tasks within 5G environments. Mobile-LLaMA supports three primary functions: IP routing analysis, packet analysis, and performance analysis.
@@ 4.2 Collection Phase
Although evidence collection or seizure traditionally involves physical tasks that require human interaction, LLMs can play a role in identifying and listing potential pieces of evidence at a crime scene. For example, in the examination of photographs or video records from a crime scene, an investigator can enlist the help of a MLLM such as LLaVa, GPT-4, or VisionLLM. These models are capable of processing information within the images and generating a text-based output, facilitating the interpretation and categorisation of visual data. Although this task may seem simple and within the capabilities of a human agent, the efficiency becomes particularly evident when dealing with a massive-scale investigation involving thousands of collected artefacts and photographs. Using an MLLM for initial processing can significantly save time, with human agents then focussing on the crucial task of verification and validation.
@@ 4.3 Preservation/Acquisition Phase
The preservation of evidence is centred on maintaining integrity. To achieve this, various tools such as EnCase and FTK Imager have been used, helping investigators streamline their work processes [155]. In the context of non-technical DF stakeholders being able to interrogate the evidence, it becomes feasible for a user to articulate their requirements/query in natural language. Subsequently, the LLM generates source code tailored to the specific need, executes the code on the data, and returns the result in consumable natural language.
LLMs specialised in code generation, such as StarCoder and Code LLaMA, can be fine-tuned and retrained for domain-specific tasks, including the preservation of disk evidence through customised code and script generation. These LLMs are capable of generating scripts or code snippets that create secure copies of disk images, metadata, and partition information, as well as automating cryptographic hashing and verification routines to maintain the evidence’s integrity through checksums. Additionally, LLMs can assist in documenting preservation steps by generating logs and summaries for each stage of the disk preservation process, thereby supporting the chain of custody during acquisitions. However, despite these capabilities, human expertise remains essential for identifying and collecting potential sources of evidence during the preservation phase, as the application of LLMs in this stage is currently limited to lower-potential tasks.
In certain instances, the gathering of live data for forensic investigations becomes crucial, particularly data collected at the crime scene. For this purpose, investigators can use DFaaS platforms such as Hansken. Hansken possesses the ability to amalgamate custom extraction APIs for data extractions, and these APIs can be developed using code-generative LLMs [20]. This approach improves the adaptability and efficiency of the investigative process.
As stated in Section 3.5, the automation of code generation and unit testing can be facilitated by autonomous agents that use LLMs as their core. AutoGen, being an open source framework, provides the means to develop AI agents tailored for specific tasks. These AutoGen agents are not only customisable and conversational, but can also operate in various modes, employing combinations of LLMs, human inputs, and various tools [129].
Automated agents, particularly those developed within frameworks such as AutoGen, can be used in the preservation phase of investigations. These agents can be assigned specific tasks, such as acquiring disk images, generating disk hashes, retrieving disk metadata, and compiling acquisition reports. By defining precise roles and tasks for AI agents, it is possible to streamline and standardise these preservation actions, improving the management of digital evidence [129, 156].
@@ 4.4 Examination Phase
This phase constitutes a crucial component of the investigation, playing a crucial role in elucidating the case through activities such as data recovery, collection, reduction, and classification. For each of these components, LLMs fine-tuned for scripting can significantly assist, especially at a larger scale. Within these components, tasks such as keyword search, file recovery, pattern matching, and fragment reassembly can be achieved with minimal technical knowledge using LLMs. LLMs can provide valuable assistance in these tasks by generating new codes, crafting regular expressions, generating passwords and/or password hash lists for decryption, and creating sample logs or files. LLMs can generate a set of instructions, queries, and Application Programming Interface (API) validations from natural language provided by a human. This opens up the possibility of integrating third-party tools like Scapy, tshark, John the Ripper, and others seamlessly into the investigative process, enhancing the toolkit available for DF investigations, and the ability to automate these processes enhances efficiency and effectiveness in the examination phase of the investigation.
The use of LAMs and VoT techniques in the examination phase can significantly enhance the efficiency of an investigation. Since LAMs and VoT specifically focus on task manipulation, investigators can offload some examination work to an LAM, which will then generate the final results from a series of subtasks. This approach can allow investigators to focus on higher-level analysis and decision making, thus streamlining the overall investigative process.
@@ 4.5 Analysis Phase
The analysis phase involves understanding the incident and obtaining a conclusive understanding based on the information collected during the examination phase. As also highlighted in Section 1.1.3, it has been demonstrated that LLMs are effective in case analysis [30]. The use of MLLMs, which possess the capability to interpret images, broadens the scope for analysing a crime case more comprehensively.
Using Gemini 1.5, Xu et al. [157] presented a tutorial on profiling a suspect’s web history through an LLM. This case study demonstrates how an LLM can help identify potential motivations, personal interests, and psychological characteristics of the suspect. In conclusion, the authors suggest that such mechanisms could power AI-assisted tools, enabling law enforcement authorities to improve the identification of cybercriminals and malicious entities.
The Digital Forensic Cybercrime Language as a Service (DFClaaS) is an innovative system developed to address the complexities of text-based cybercrime [158]. Using natural NLP techniques, including LLMs, sentiment analysis, and lexicon analysis, DFClaaS aims to improve capabilities in incident reporting, analysis, and investigation. The primary objectives of DFClaaS include implementing microservices to address specific challenges, proposing an advanced system to improve incident handling, and providing valuable tools for DF investigators. Designed to serve individual users, organisations, and forensic professionals, DFClaaS is a versatile and effective resource in the ongoing fight against cybercrime.
LLMs can be specifically fine-tuned for the analysis of various data types, including log files, email contents, chat transcripts, call records, file metadata, hex dumps, memory dumps, and registry hives. Incorporating contents such as event logs, timestamps, and network traffic captures further enables the effective recreation of incidents by correlating each data set with the assistance of LLMs. In addition, MLLMs that are audio and video specific can assist in analysing content within these formats. This specialised capability can significantly reduce the time investigators spend analysing audio and video data during investigations.
The use of automated agents can effectively distribute the analysis workload. Moreover, leveraging Augmented Large Language Models (ALLMs) and RAG techniques can improve knowledge retrieval in real time, thus improving the accuracy of analysis and decision-making processes [159]. For example, integrating a source of intelligence with an RAG system can assist investigators in connecting the dots during a DF investigation.
Other than these applications, LLMs can increase productivity through enhanced information correlation during the analysis phase. Shafee et al. [160] suggest that LLMs hold significant potential for data correlation and cybersecurity applications. The referenced study evaluated the performance of various LLM-based chatbots, including ChatGPT, GPT4all, Dolly, Stanford Alpaca, Alpaca-LoRA, Falcon, and Vicuna, specifically for text classification and Named Entity Recognition (NER) tasks using OSINT data. The findings indicated that, although the commercial chatbot GPT-4 and the open-source GPT4all performed well in text classification, all tested LLM-based chatbots showed limitations and were less effective for cybersecurity entity recognition compared to specialised models. The study concludes that there remains room for improvement.
@@ 4.6 Reporting Phase
The quality and validity of the evidence, along with the thoroughness of the analysis, are encapsulated in the final report. The reporting phase holds significant weight, as the entire judgement may hinge on this crucial stage. Notably, DF is experiencing heightened scrutiny about the quality of the reports, emphasising the importance of precision and clarity in this phase [161]. As pointed out by Champod et al. [162], there is no standard framework for evaluating and reporting scientific findings to authorities and stakeholders. To provide assistance and alleviate scrutiny, incorporating LLMs for report creation is a viable solution.
While LLMs are inherently non-deterministic, adhering to investigation standards such as ISO/IEC 27043:2015 can establish robust processes around data integrity and evidence handling, even though these standards do not directly address the randomness or variability in LLM outputs. The ISO/IEC 27043:2015 standard provides guidelines for a consistent DF investigation framework, focusing on maintaining procedural rigour rather than modifying model behaviour. Although it does not directly resolve issues of LLM determinism, it can serve as a protocol to ensure that procedures involving LLMs uphold investigative standards and maintain integrity throughout the process [163].
A preliminary feasibility study by Michelet and Breitinger [164] highlighted the potential of LLMs to assist in automating forensic report generation. These models can facilitate the creation of structured sections, such as methodologies, data analysis, and summaries, by generating coherent, case-specific insights from forensic data. Additionally, LLMs could automate the production of reports in alternative formats, such as HTML or , which are frequently used for dynamic, web-based, or highly technical documentation.
@@
4.7 Other Possibilities
Scanlon et al. [13] highlights that LLMs can play an important role in teaching scenarios. This involvement extends to activities such as storyboarding, creation of synthetic content, and synthetic character profiling.
Fine-tuned models could further enhance training by generating more complex, realistic case examples that challenge trainees with nuanced scenarios, providing a robust foundation for practical skills development. These models may also help translate technical findings into accessible language, facilitating communication of insights to non-specialists, such as judges or other stakeholders.
4.8 Discussion on Potential for LLMs in DF
To provide a comprehensive understanding of the potential use of LLMs, Table 3 clarifies the sample functionalities within the framework of the National Institute of Standards and Technology (NIST) Computer Forensics Tool Testing Program (CFTT), highlighting the usability of LLMs and example prompts. The CFTT project establishes overarching specifications to assess the capabilities of tools, a framework adopted by numerous prominent free and commercial tools  [^5] https://www.nist.gov/itl/ssd/software-quality-group/computer-forensics-tool-testing-program-cftt.
The potential for having a positive impact on the typical phases of the investigation increases as one progresses through the typical order of the phases. For example, there is little improvement that can be made by an LLM or automated scripting during the identification or acquisition phases, but significant potential for aiding investigators in the reporting phase [164, 156] – these are first and foremost large language models. The low/medium/high potential outlined below evaluates each DF phase based on three key requirements: reliance on human expertise, physical versus digital evidence handling, and scope for automation, as explained below.
  • 1.
    Low Potential for the Identification and Collection Phases
    • (a)
      High dependency on human involvement, expertise, and/or specialised knowledge.
    • (b)
      Involves extensive handling of physical evidence.
    • (c)
      Limited or no feasibility for automation.
  • 2.
    Medium Potential for the Preservation Phase
    • (a)
      Requires some level of human involvement or expertise, but is not critical to the process.
    • (b)
      Primarily deals with digital evidence, with minimal physical evidence handling.
    • (c)
      Feasible for automation to a significant extent.
  • 3.
    High Potential for the Examination, Analysis and Reporting Phases
    • (a)
      Human involvement is needed for expert verification of the conducted analysis.
    • (b)
      Exclusively focused on digital evidence.
    • (c)
      Many common tasks are suitable for significant support from LLMs.
With these possibilities, the scope for research in DF is vast. Future research could be extended to the generation of digital forensic reports, as well as the summarisation of these reports for non-technical users. This would save time, but can also lead to more consistent documentation compared to manual documentation. Given the capacity of LLMs to manage large textual datasets, exploring pattern recognition holds significant value, particularly for investigations requiring the detection of anomalies or outliers in chats, log events, or emails.
In addition, LLMs’ ability to interpret the tone of messages or chats enables their application in the sentiment analysis of text-based evidence. There is also potential in fine-tuning LLMs for domain-specific tasks, such as network forensics, where LLMs could analyse log files and application data related to specific activities. Automating LLM-based DF tools could further enable investigators to generate customised reports using natural language queries.
A critical future research direction lies in the ethical and legal considerations of LLM-generated content. As the application of LLMs is still emerging, future studies should focus on developing appropriate benchmarks, standardisation protocols, and addressing legal aspects to ensure responsible use of this technology.